# Non-convergence of the Navier-Stokes equations toward the Euler equations in weak Besov spaces

### v1

@@ Abstract
In this paper, we consider the inviscid limit problem to the higher dimensional incompressible Navier-Stokes equations in the whole space. It was proved in [16, J. Funct. Anal., 276 (2019)] that given initial data $u_{0}\in B^{s}_{p,r}$ with $1\leq r<\infty$, the solutions of the Navier-Stokes equations converge strongly in $B^{s}_{p,r}$ to the Euler equations as the viscosity parameter tends to zero. In the case when $r=\infty$, we prove the failure of the $B^{s}_{p,\infty}$-convergence of the Navier-Stokes equations toward the Euler equations in the inviscid limit.
@@ 1 Introduction
The motion of an incompressible viscous fluid with constant density is governed by the following Navier-Stokes equations:
@@
When the viscocity vanishes ($\varepsilon=0$), then the Navier-Stokes equations (NS) reduces to the Euler equations for ideal incompressible fluid
@@
A classical problem in fluid mechanics is the approximation in the limit $\varepsilon\to 0$ of vanishing viscosity (also called inviscid limit or zero-viscosity limit) of solutions of the Euler equations by solutions of the incompressible Navier-Stokes equations, which is naturally associated with the physical phenomena of turbulence and of boundary layers. The vanishing viscosity problem is closely related to uniqueness of solutions to the Euler equations, because the methods used to prove uniqueness can often be applied to show vanishing viscosity. In the presence of boundaries, the zero-viscosity limit problem is more complicated and essentially difficult as it is closely related to the validity of the Prandtl equation for the formation of boundary layers. We refer to see the recent results by Maekawa [20], Constantin, Kukavica and Vicol [10, 11], Gérard et al. [15] and the references therein. For initial data that is analytic only close to the boundary of the domain, Kukavica, Vicol and Wang [26, 27] proved that the solution of the Navier-Stokes equations converges in the vanishing viscosity limit to the solution of the Euler equations.
@@
Most of the progress has been made in the two-dimensional case without boundary (assuming the domain to be the whole plane $\mathbb{R}^{2}$ of the torus $\mathbb{T}^{2}$) since both vortex stretching and boundary effects are absent. In particular, the problem of the convergence of smooth viscous solutions of the Navier-Stokes equations to the Eulerian one as $\varepsilon\to 0$ is well understood and has been studied in many literatures, see for example [9, 29, 19]. Majda [21] showed that under the assumption $u_{0}\in H^{s}$ with $s>\frac{d}{2}+2$, the solutions to (NS) converge in $L^{2}$-norm as $\varepsilon\to 0$ to the unique solution to (E) and the convergence rate is of order $(\varepsilon t)^{\frac{1}{2}}$. Masmoudi [23] improved the result and proved the convergence in $H^{s}$-norm under the weaker assumption $u_{0}\in H^{s}$ with $s>\frac{d}{2}+1$. Next we shall recall the consideration of Yudovich solutions. Yudovich [31] established the uniqueness of solutions to the Euler equations in the space $C\left(\mathbb{R}^{+};L^{2}\left(\mathbb{R}^{2}\right)\right)\times L_{\text{loc }}^{\infty}\left(\mathbb{R}^{+};L^{2}\left(\mathbb{R}^{2}\right)\right)$ assuming that initial data lies to $L^{2}\left(\mathbb{R}^{2}\right)$ and initial vorticity belongs to $L^{p}\left(\mathbb{R}^{2}\right)\cap L^{\infty}\left(\mathbb{R}^{2}\right)$ for some $p<\infty$. For this uniqueness class, Chemin [5] showed that solutions of the Navier-Stokes equations with initial vorticity lies in $L^{2}(\mathbb{R}^{2})\cap L^{\infty}(\mathbb{R}^{2})$ converge in the zero-viscosity limit to a solution of the Euler equations. Kelliher [25] resolved the same inviscid limit as Chemin [5] with the assumption that the vorticity is unbounded. We also mention that, in the case of that initial velocity has finite energy and initial vorticity is bounded in the plane, Cozzi [12, 13] showed that the unique solutions of the Navier-Stokes equations converge to the unique solution of the Euler equations in the $L^{\infty}$-norm uniformly over finite time as viscosity approaches zero. In the case of two dimensional (torus or the whole space) and rough initial data, by taking greater advantage of vorticity formulation, more beautiful results were obtained quantitatively (see for example [2, 14, 6, 8, 25, 32] and the references therein).
@@
(1)  Uniform bounds: there exists $C=C(s,p,r,d)>0$ such that
@@
(2) Inviscid limit: $\lim\limits_{\varepsilon\downarrow 0}\left\|u^{\rm NS}_{\varepsilon}(t,u_{0})-u^{\rm E}(t,u_{0})\right\|_{L_{T}^{\infty}B^{s}_{p,r}}=0,$ namely, for any $u_{0}\in B^{s}_{p,r}$, $\forall\eta>0$, $\exists\ \varepsilon_{0}=\varepsilon_{0}(u_{0},\eta,T)$, $\forall\varepsilon\leq\varepsilon_{0}$, such that
@@
However, we should be noted that it left an open problem for the end-point case $r=\infty$ in the inviscid limit of the Navie-Stokes equations in Besov spaces. Precisely speaking, the statement (2) is a consequence of the strong convergence of the solutions to (NS) with the given initial data belonging in $B^{s}_{p,r}$ with $1\leq r<\infty$. An interesting Question appears:
For any $u_{0}\in B^{s}_{p,\infty}$, in the inviscid limit $\varepsilon\to 0$, whether or not the $B^{s}_{p,\infty}$-convergence
@@
In this paper, we will give a negative answer: there can be no strong convergence in $B^{s}_{p,\infty}$, i.e.,
@@
Now we state main result.
@@
For any $u_{0}\in U_{R}\subseteq B^{s}_{p,\infty}(\mathbb{R}^{d})$, then a family of solutions $\left\{u^{\rm NS}_{\varepsilon}(t,u_{0})\right\}_{\varepsilon>0}$ to (NS) do not converge to the solution $u^{\rm E}(t,u_{0})$ of (E) in $B^{s}_{p,\infty}(\mathbb{R}^{d})$. More precisely, there exists a sequence initial data $u_{0}\in U_{R}$ such that for a short time $T_{0}\leq T$
$\liminf_{\varepsilon\downarrow 0}\left\|u^{\rm NS}_{\varepsilon}(t,u_{0})-u^{\rm E}(t,u_{0})\right\|_{L^{\infty}_{T_{0}}B^{s}_{p,\infty}}\geq\eta_{0},$
with some positive constant $\eta_{0}$ depending on $d$, but independent of $\varepsilon$.
@@
To the best of our knowledge, Theorem 1.2 is new. In fact, it was proved in [16] that given initial data $u_{0}\in B^{s}_{p,r}$ with $1\leq r<\infty$ and for some $T>0$, the solutions of the Navier-Stokes equations converge strongly in $L^{\infty}_{T}B^{s}_{p,r}$ to the Euler equations as the viscosity parameter tends to zero while our Theorem 1.2 shows the non-convergence of the solutions for the Navier-Stokes equations toward the Euler equations in the inviscid limit under the framework of weak Besov spaces $B^{s}_{p,\infty}$.
@@
Due to the incompressible condition, the pressure can be eliminated from (NS) and (E). In fact, applying the Leray operator $\mathcal{P}$ to (NS) and (E), respectively, then we have
@@
$u^{\rm NS}_{\varepsilon}(t,u_{0})=e^{t\varepsilon\Delta}u_{0}-\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}(u^{\rm NS}_{\varepsilon}\cdot\nabla u^{\rm NS}_{\varepsilon})\mathrm{d}\tau,$ (1.2)
@@
Step 1: At a first sight, to ensure that the difference between the solutions $u^{\rm NS}_{\varepsilon}$ to (NS) and $u^{\rm E}$ to (E) in the $B^{s}_{p,\infty}$-topology is bounded below by a positive constant at any later time, we need to construct the initial data sequence $u_{0}\in B^{s}_{p,\infty}$ to satisfy that the difference between the solutions $e^{t\varepsilon\Delta}u_{0}$ and $u_{0}$ in the $B^{s}_{p,\infty}$-topology is bounded below by a positive constant at any later time. In order to see this, let us introduce the following toy-model:
A toy-model: 1-D heat equation. Let $\varepsilon_{n}:=2^{-2n}$ with $n\gg 1$. The Cauchy problem
@@
Let $\phi$ be given in subsection 3.1 and set $u_{0}(x)=2^{-ns}\phi(x)\cos\left(\frac{17}{12}2^{n}x\right)$, then $u^{\varepsilon_{n}}\nrightarrow u_{0}$ in $B^{s}_{p,\infty}(\mathbb{R})$. In fact, for $n\gg 1$, one has for $t>0$
$\left\|e^{\varepsilon_{n}t\partial^{2}_{xx}}u_{0}-u_{0}\right\|_{B^{s}_{p,\infty}}\geq\left(e^{9t/16}-1\right)\left\|\phi(x)\cos\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{p}}\nrightarrow 0\quad\text{as}\;n\to\infty.$
@@
is smaller than the leading term $e^{t\varepsilon\Delta}u_{0}-u_{0}$. We can decompose the above error term into four terms (see (3.4) in subsection 3.4) and estimate them by subtle argument (see Propositions 3.3-3.4 in subsection 3.3).
@@ 2 Preliminaries
  • Let us recall that for all $u\in\mathcal{S}^{\prime}$, the Fourier transform $\mathcal{F}u$, also denoted by $\widehat{u}$, is defined by
    $\mathcal{F}u(\xi)=\widehat{u}(\xi)=\int_{\mathbb{R}^{d}}e^{-\mathrm{i}x\cdot\xi}u(x)\mathrm{d}x\quad\text{for any}\;\xi\in\mathbb{R}^{d}.$
  • The inverse Fourier transform allows us to recover $u$ from $\widehat{u}$:
    $u(x)=\mathcal{F}^{-1}\widehat{u}(x)=(2\pi)^{-d}\int_{\mathbb{R}^{d}}e^{\mathrm{i}x\cdot\xi}\widehat{u}(\xi)\mathrm{d}\xi.$
@@
    In $\mathbb{R}^{d}$, $\mathcal{P}$ can be defined by $\mathcal{P}=\mathrm{Id}+(-\Delta)^{-1}\nabla{\rm{div}}$, or equivalently, $\mathcal{P}=(\mathcal{P}_{ij})_{1\leqslant i,j\leqslant d}$, where $\mathcal{P}_{ij}\equiv\delta_{i,j}+R_{i}R_{j}$ with $\delta_{i,j}$ being the Kronecker delta ($\delta_{i,j}=0$ for $i\neq j$ and $\delta_{i,i}=0$) and $R_{i}$ being the Riesz transform with symbol $-\mathrm{i}\xi_{1}/|\xi|$. Obviously, $\mathcal{Q}=-(-\Delta)^{-1}\nabla{\rm{div}}$, and if $\mathord{{\rm div}}\,u=\mathord{{\rm div}}\,v=0$, it holds that $\mathcal{Q}(u\cdot\nabla v)=\mathcal{Q}(v\cdot\nabla u).$
@@
$\Delta_{j}u=\varphi(2^{-j}D)u,\;\;\text{if}\;j\in\mathbb{Z},$
@@
Due to the fact $\varphi(\xi)\equiv 1$ for $\frac{4}{3}\leq|\xi|\leq\frac{3}{2}$, for $-1\leq j,n\in\mathbb{Z}$, we have
@@
Before proving Theorem 1.2, we recall the well-posedness result. Assume $(s,p)$ satisfies the condition (1.1). It is well-known, by the classical well-posedness theory, that for fixed $\varepsilon>0$, there exists $T=T(\|u_{0}\|_{B^{s}_{p,\infty}})>0$ independent of $\varepsilon$ such that (NS) has a unique solution $u$ in $L^{\infty}([0,T];B^{s}_{p,\infty})$.

### v2

# Non-convergence of the Navier-Stokes equations toward the Euler equations in the endpoint Besov spaces

@@ Abstract
In this paper, we consider the inviscid limit problem to the higher dimensional incompressible Navier-Stokes equations in the whole space. It was proved in [16, J. Funct. Anal., 276 (2019)] that given initial data $u_{0}\in B^{s}_{p,r}$ with $1\leq r<\infty$, the solution of the Navier-Stokes equations converges strongly in $B^{s}_{p,r}$ to the solution of the Euler equations as the viscosity parameter tends to zero. In the case when $r=\infty$, we prove the failure of the $B^{s}_{p,\infty}$-convergence of the Navier-Stokes equations toward the Euler equations in the inviscid limit.
@@ 1 Introduction
The motion of an incompressible viscous fluid with constant density is governed by the following Navier-Stokes equations
@@
When the viscocity vanishes ($\varepsilon=0$), the Navier-Stokes equations (NS) reduces to the Euler equations for ideal incompressible fluid
@@
A classical problem in fluid mechanics is the approximation in the limit $\varepsilon\to 0$ of vanishing viscosity (also called inviscid limit or zero-viscosity limit) of solution of the Euler equations by solution of the incompressible Navier-Stokes equations, which is naturally associated with the physical phenomena of turbulence and of boundary layers. The vanishing viscosity problem is closely related to uniqueness of solution to the Euler equations, because the methods used to prove uniqueness can often be applied to show vanishing viscosity. One of the most important uniqueness results in $\mathbb{R}^{2}$ is due to Yudovich [30] and the vanishing viscosity results is due to Chemin [5]. As mentioned in Kelliher [25], given the uniqueness of solution to the Euler equations, the compactness argument would imply the vanishing viscosity limit. In the presence of boundaries, the zero-viscosity limit problem is more complicated and essentially difficult as it is closely related to the validity of the Prandtl equation for the formation of boundary layers. We refer to see the recent results by Maekawa [20], Constantin, Kukavica and Vicol [10, 11], Gérard et al. [15] and the references therein. For initial data that is analytic only close to the boundary of the domain, Kukavica, Vicol and Wang [26, 27] proved that the solution of the Navier-Stokes equations converges in the vanishing viscosity limit to the solution of the Euler equations.
@@
Most of the progress has been made in the two-dimensional case without boundary (assuming the domain to be the whole plane $\mathbb{R}^{2}$ of the torus $\mathbb{T}^{2}$) since both vortex stretching and boundary effects are absent. In particular, the problem of the convergence of smooth viscous solution of the Navier-Stokes equations to the Eulerian one as $\varepsilon\to 0$ is well understood and has been studied in many literatures, see for example [9, 28, 19]. Majda [21] showed that under the assumption $u_{0}\in H^{s}$ with $s>\frac{d}{2}+2$, the solution to (NS) converges in $L^{2}$-norm as $\varepsilon\to 0$ to the unique solution to (E) and the convergence rate is of order $(\varepsilon t)^{\frac{1}{2}}$. Masmoudi [23] improved the result and proved the convergence in $H^{s}$-norm under the weaker assumption $u_{0}\in H^{s}$ with $s>\frac{d}{2}+1$. Next we shall recall the consideration of Yudovich solution. Yudovich [30] established the uniqueness of solution to the Euler equations in the space $C\left(\mathbb{R}^{+};L^{2}(\mathbb{R}^{2})\right)\times L_{\text{loc }}^{\infty}\left(\mathbb{R}^{+};L^{2}(\mathbb{R}^{2})\right)$ assuming that initial data lies to $L^{2}(\mathbb{R}^{2})$ and initial vorticity belongs to $L^{p}(\mathbb{R}^{2})\cap L^{\infty}(\mathbb{R}^{2})$ for some $p<\infty$. For this uniqueness class, Chemin [5] showed that solution of the Navier-Stokes equations with initial vorticity liying in $L^{2}(\mathbb{R}^{2})\cap L^{\infty}(\mathbb{R}^{2})$ converges in the zero-viscosity limit to a solution of the Euler equations. Kelliher [25] resolved the same inviscid limit as Chemin [5] with the assumption that the vorticity is unbounded. We also mention that, in the case of that initial velocity has finite energy and initial vorticity is bounded in the plane, Cozzi [12, 13] showed that the unique solution of the Navier-Stokes equations converges to the unique solution of the Euler equations in the $L^{\infty}$-norm uniformly over finite time as viscosity approaches zero. In the case of two dimensional (torus or the whole space) and rough initial data, by taking greater advantage of vorticity formulation, more beautiful results were obtained quantitatively (see [2, 14, 6, 8, 25, 31] and the references therein).
@@
(1) (Uniform bounds): there exists $C=C(s,p,r,d)>0$ such that
@@
(2) (Inviscid limit): $\lim\limits_{\varepsilon\downarrow 0}\left\|u^{\rm NS}_{\varepsilon}(t,u_{0})-u^{\rm E}(t,u_{0})\right\|_{L_{T}^{\infty}B^{s}_{p,r}}=0,$ namely, for any $u_{0}\in B^{s}_{p,r}$, $\forall\eta>0$, $\exists\ \varepsilon_{0}=\varepsilon_{0}(u_{0},\eta,T)$, $\forall\varepsilon\leq\varepsilon_{0}$, such that
@@
However, it should be noted that we left an open problem for the case $r=\infty$ in the inviscid limit of the Navie-Stokes equations in the endpoint Besov spaces $B^{s}_{p,\infty}$. Notice that again, the statement (2) is a consequence of the strong convergence of the solution to (NS) with the given initial data belonging in $B^{s}_{p,r}$ with $1\leq r<\infty$. An interesting Question appears:
Given initial data $u_{0}\in B^{s}_{p,\infty}$, in the inviscid limit $\varepsilon\to 0$, whether or not the $B^{s}_{p,\infty}$-convergence
@@
In this paper, we will give a negative answer: By constructing a particular initial data $u_{0}\in B^{s}_{p,\infty}$, we find that
@@
Now we state the main result.
@@
Given initial data $u_{0}\in U_{R}\subseteq B^{s}_{p,\infty}(\mathbb{R}^{d})$, then
(1) there exists a positive time $T_{1}$ such that (E) has a unique solution $u^{\rm E}(t,u_{0})\in L^{\infty}_{T_{1}}B^{s}_{p,\infty}$;
(2) for all $\varepsilon>0$, there exists a positive time $T_{2}\geq T_{1}$ such that (NS) has a unique solution $u^{\rm NS}_{\varepsilon}(t,u_{0})\in L^{\infty}_{T_{2}}B^{s}_{p,\infty}$;
(Non-convergence): there exists an initial data $u_{0}\in B^{s}_{p,\infty}(\mathbb{R}^{d})$ such that the solution $u^{\rm NS}_{\varepsilon}(t,u_{0})$ of (NS) does not converge to the solution $u^{\rm E}(t,u_{0})$ of (E) for small $t\in(0,T_{1}]$ in $B^{s}_{p,\infty}(\mathbb{R}^{d})$ as $\varepsilon\downarrow 0$. More precisely, for all $\varepsilon>0$, there exists an initial data $u_{0}\in B^{s}_{p,\infty}(\mathbb{R}^{d})$ and a small enough $\delta>0$ such that for a short time $t=\delta\varepsilon$
$\left\|u^{\rm NS}_{\varepsilon}(t,u_{0})-u^{\rm E}(t,u_{0})\right\|_{B^{s}_{p,\infty}}\geq\eta_{0},$ (1.2)
with some positive constant $\eta_{0}$ depending on $d$ and $\delta$ but independent of $\varepsilon$.
@@
The novelty of Theorem 1.2 is part (3), while part (1) and (2) is classical (see [16, Remark 2]).
@@
Due to the incompressibility condition, the pressure can be eliminated from (NS) and (E). In fact, applying the Leray operator $\mathcal{P}$ to (NS) and (E), respectively, then we have
@@
$u^{\rm NS}_{\varepsilon}(t,u_{0})=e^{t\varepsilon\Delta}u_{0}-\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}\left(u^{\rm NS}_{\varepsilon}\cdot\nabla u^{\rm NS}_{\varepsilon}\right)\mathrm{d}\tau,$ (1.3)
@@
Step 1: At a first sight, to ensure that the difference between the solution $u^{\rm NS}_{\varepsilon}$ to (NS) and $u^{\rm E}$ to (E) in the $B^{s}_{p,\infty}$-topology is bounded below by a positive constant at any later time, we need to construct special initial data $u_{0}\in B^{s}_{p,\infty}$ to satisfy that the difference between $e^{t\varepsilon\Delta}u_{0}$ and $u_{0}$ in the $B^{s}_{p,\infty}$-topology is bounded below by a positive constant at any later time. In order to see this, let us introduce the following toy-model:
A toy-model: 1-D heat equation. Assume that $(s,p)$ satisfies the condition (1.1) with $d=1$. Let $\varepsilon_{n}:=2^{-2n}$ with $n\gg 1$. The Cauchy problem
@@
Let $\phi$ be given in subsection 3.1 and set $u_{0}(x)=2^{-ns}\phi(x)\cos\left(\frac{17}{12}2^{n}x\right)$, then $u^{\varepsilon_{n}}-u_{0}\nrightarrow 0$ in $B^{s}_{p,\infty}(\mathbb{R})$. In fact, for $n\gg 1$, notice that the frequency of $u_{0}$ is supported in the annulus $\{\xi\in\mathbb{R}:\ \frac{33}{24}2^{n}\leq|\xi|\leq\frac{35}{24}2^{n}\}$, one has for small $t>0$ and for some positive constants $c_{1},C_{1},C_{2}$ independent of $t$ and $n$
$C_{2}t\geq$ $~\left\|e^{\varepsilon_{n}t\partial^{2}_{xx}}u_{0}-u_{0}\right\|_{B^{s}_{p,\infty}}=2^{ns}\big{\|}\Delta_{n}\big{(}e^{t\varepsilon_{n}\partial^{2}_{xx}}u_{0}-u_{0}\big{)}\big{\|}_{L^{p}}$
$\geq$ $~t2^{ns}\big{\|}{\Delta}_{n}\varepsilon_{n}\partial^{2}_{xx}u_{0}\big{\|}_{L^{p}}-\int_{0}^{t}2^{ns}\left\|{\Delta}_{n}\varepsilon_{n}\partial^{2}_{xx}\big{(}e^{\tau\varepsilon_{n}\partial^{2}_{xx}}u_{0}-u_{0}\big{)}\right\|_{L^{p}}\mathrm{d}\tau$
$\geq$ $~ct\left\|\phi(x)\cos\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{p}}-C\int_{0}^{t}2^{ns}\left\|{\Delta}_{n}\big{(}e^{\tau\varepsilon_{n}\partial^{2}_{xx}}u_{0}-u_{0}\big{)}\right\|_{L^{p}}\mathrm{d}\tau$
$\geq$ $~c_{1}t-C_{1}t^{2}.$
For more details see subsection 3.1.
@@
is smaller than the leading term $e^{t\varepsilon\Delta}u_{0}-u_{0}$. We can decompose the above error term into four terms (see (3.12) in subsection 3.4) and estimate them by some basic analytical techniques (see Propositions 3.3-3.4 in subsection 3.3).
@@ 2 Preliminaries
  • For all $u\in\mathcal{S}(\mathbb{R}^{d})$ (Schwartz space), the Fourier transform $\mathcal{F}u$ (also denoted by $\widehat{u}$) is defined by
    $\left(\mathcal{F}u\right)(\xi)=\widehat{u}(\xi)=\int_{\mathbb{R}^{d}}e^{-\mathrm{i}x\cdot\xi}u(x)\mathrm{d}x\quad\text{for any}\;\xi\in\mathbb{R}^{d}.$
  • For all $u\in\mathcal{S}(\mathbb{R}^{d})$, the inverse Fourier transform is defined by $(\mathcal{F}^{-1}u)=(2\pi)^{-d}\left(\mathcal{F}u\right)(-\xi)$, namely,
    $(\mathcal{F}^{-1}u)(x)=(2\pi)^{-d}\int_{\mathbb{R}^{d}}e^{\mathrm{i}x\cdot\xi}u(\xi)\mathrm{d}\xi.$
  • For all $F\in\mathcal{S^{\prime}}(\mathbb{R}^{d})$ (space of tempered distributions), the Fourier transform $\mathcal{F}F=\widehat{F}$ is defined by
    $\langle\mathcal{F}F,u\rangle=\langle F,\mathcal{F}u\rangle\quad\text{for all}\;u\in\mathcal{S}(\mathbb{R}^{d}).$
@@
    In $\mathbb{R}^{d}$, $\mathcal{P}$ can be defined by $\mathcal{P}=\mathrm{Id}+(-\Delta)^{-1}\nabla{\rm{div}}$, or equivalently, $\mathcal{P}=(\mathcal{P}_{ij})_{1\leqslant i,j\leqslant d}$, where $\mathcal{P}_{ij}\equiv\delta_{i,j}+R_{i}R_{j}$ with $\delta_{i,j}$ being the Kronecker delta ($\delta_{i,j}=0$ for $i\neq j$ and $\delta_{i,i}=0$) and $R_{i}$ being the Riesz transform with symbol $\mathrm{i}\xi_{1}/|\xi|$.
@@
$\dot{\Delta}_{j}u=\varphi(2^{-j}D)u,\;\;\text{if}\;j\in\mathbb{Z},$
@@
[…]